OpenStudy (anonymous):

please help.? gingers baby weighed on 3 3/8 pounds when it was born. the doctor said it could not leave until it weighed 5 1/2 pounds. how many pounds must the baby gain before it can go home?

OpenStudy (anonymous):

5 1/2 - 3 3/8 = 11/2 -27/8 = 44/8 -27/8 = 17/8 pounds